Database last updated on Mon, 02 Feb 2026 06:40:53 +0000 by localhost

Instrument constant changes for the week before 2024022 (2 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 16 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
02620240222326.9

Non-critical warnings from b-files' comments for days between 2024022 and 2024022 (0 sec, 6 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
dp:Doing dispersion with internal HG.199B02224.026TORON#02611:49:461-1
skc:HOME key pressed299B02224.026TORON#02612:34:151-1
dsp:Doing dispersion test with Hg (UV port)299B02224.026TORON#02613:23:421-1
checkhg:Last HG had to adjust the micrometer by at least 1199B02224.026TORON#02614:51:141-1
hg:DEL key pressed199B02224.026TORON#02615:47:231-1
b1:Going to menu199B02224.026TORON#02615:50:331-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhgrssluv
022TORONTO#026HOME key pressed,Must do HG now,sch026,Waiting until ZA =,Waiting until ZA = 170 now 124.1,Waiting until ZA = 170 now 125.9,Waiting until ZA = 170 now 127.6,Waiting until ZA = 170 now 129.4,Waiting until ZA = 170 now 131.2,Waiting until ZA = 170 now 133,Waiting until ZA = 170 now 134.7,Waiting until ZA = 170 now 136.4,Waiting until ZA = 170 now 138.1,Waiting until ZA = 170 now 139.8,Waiting until ZA = 170 now 141.4,Waiting until ZA = 170 now 143,Waiting until ZA = 170 now 144.5,Waiting until ZA = 170 now 146,Waiting until ZA = 170 now 147.4,Waiting until ZA = 170 now 148.8,Waiting until ZA = 170 now 150.1,Waiting until ZA = 170 now 151.2,Waiting until ZA = 170 now 152.3,Waiting until ZA = 170 now 153.3,Waiting until ZA = 170 now 154.1,Waiting until ZA = 170 now 154.7,Waiting until ZA = 170 now 155.2,Waiting until ZA = 170 now 155.6,Waiting until ZA = 170 now 155.7,Waiting until ZA = 95 now 94.322210022171
(0 sec)

Dead time (1 sec)

File nameCountSerialExpected (+/-5) Measured 30-day history
B02224.02622#0263129